OpenStudy (anonymous):

Help me solve the quadratic formula? 2x^2 = 200 See the thing about this question is that, the answer is quite obvious. 2 x 10 squared equals 100, and it's simple as that if you just think about it. Would it honestly be necessary to work out a whole quadratic formula? To explain what they're asking for, their showing me an an example of how a student wrote their formula down, but they think it's too long and want me to give a simple, more better method? The method I eventually worked up was; x = -10 pm sqrt{(0)^2 - 4(2)(-200)} x = 0 pm sqrt{1600} x = pm \frac{ 40 }{ 4 }

OpenStudy (anonymous):

\[x = -0 \pm \sqrt{(0)^2 - 4(20)(-200)}\] \[x = 0 \pm \sqrt{1600}\] \[x = \pm \frac{ 40 }{ 4 }\] x = 10 and - 10 {10, -10} Is there an easier method than what I stated in my first post?

OpenStudy (solomonzelman):

divide both sides by 2. take the square root of both sides.

OpenStudy (anonymous):

the solution(s) would be x=-10 or x=10.